OpenStudy (morgengarner):

Simplify the expression. The quantity m raised to the power of negative 6 multiplied by n raised to the power of negative 3 end of quantity divided by the quantity m raised to the power of negative 13 multiplied by n raised to the power of negative 1 end of quantity. A. m3 n12 B. N raised to the power of negative 9 divided by n raised to the power of negative 14. C. M raised to the power of 7 divided by n raised to the power of 2. D. m7 n2

OpenStudy (calculusxy):

I suggest for you to actually write it down with numbers and operating symbols. \[\large \frac{m^{-6} \times n^{-3}}{m^{-13} \times n^{-1}}\]

OpenStudy (calculusxy):

When you are dividing exponents with the same base, you have to subtract the exponents. \[\large \frac{x^a}{x^b} = x^{a-b}\]

OpenStudy (calculusxy):

And if you have negative exponent you need to make sure that they turn positive by changing the place (numerator or denominator) of the base. That base will then have the same exponent but now it will be positive. \[\large x^{-a} = \frac{1}{x^a}\] or \[\large \frac{1}{x^{-1}} = \frac{x^1}{1} = x^1\]

OpenStudy (calculusxy):

So for your problem do: \[\large m^{-6 - (-13)}\] and \(\large n^{-3 - (-1)}\)

OpenStudy (morgengarner):

thank you

OpenStudy (calculusxy):

No problem